Two stock investments cost $15 000. One stock then had a 40% gain and the other a 10% loss. If the net profit is $2000, how much was invested in each stock???
Answer by KnightOwlTutor(293)

X =Stock 1
15,000-X=Stock 2
Stock 1 had a 40% gain convert to decimal .40
Stock 2 had a 10% loss -10% convert to decimal -.10
(profit or loss of Stock 1)+(profit or loss of Stock 2)=net profit
.40X-.10(15000-X)=2000
Use the distributive property
.40X-1500+.10X=2000
remember (-)(-)=(+)
.5X-1500=2000
add 1500 to both sides
.5X=3500
divide .5 on both sides
X=$7000= Amount invested in Stock 1
15,000-7000=$8000 Amount invested in Stock 2